OpenStudy (anonymous):

solve 2x-7=6x-15 A.2 B.3 C.6 D.5

OpenStudy (arabpride):

combine like terms: 2x-7=6x-15 -2x -2x -7 = 4x - 15 +15 +15 8 = 4x /4 /4 2=x x=2

OpenStudy (arabpride):

@chelsea49

OpenStudy (anonymous):

yes

OpenStudy (arabpride):

You're Welcome *^.^*

OpenStudy (anonymous):

:)

OpenStudy (timaashorty):

just plug in all your choices. \[2(2)- 7 = 6(2) - 15 \] \[4-7 = 12-15 \] \[-3 = -3 \] therefore, \[\huge \ it \ is \ 2 \]

OpenStudy (arabpride):

Yerp *^.^*

OpenStudy (anonymous):

i need help again

OpenStudy (timaashorty):

Post it in another question (:

OpenStudy (arabpride):

Yerp ~ different post after this one nd we'll be glad to help ya out ~.^

OpenStudy (anonymous):

solve [x]+3=7 A. {4} B. {-4,4} C. {-10,-4} D. {-10}

OpenStudy (arabpride):

can u post the question on a different post? that way both of us get good out of it ~ it raises our SS ~.^ ill help ya NP

OpenStudy (anonymous):

ok